# Graph-Based Approximate Nearest Neighbor Search Revisited: Theoretical Analysis and Optimization

### v2

@@ Appendix F Complexity Analysis Details
Below is a detailed complexity analysis of each component in the Vamana algorithm. Complexity of GreedySearch. The GreedySearch algorithm iteratively selects the closest neighbor to the query and follows a path to an approximate nearest neighbor, evaluating all neighbors at each step. According to [12], the complexity of GreedySearch is determined by the product of the average path length and the average degree. Theorem 2 establishes that the path length is $O(\log n)$. Thus, for a graph with degree $K$, the search complexity is $O(K\cdot\log n)$.
@@
Complexity of Pruning. In the SNG pruning algorithm, the outer loop checks if the candidate set $S$ is empty. If not, it adds an edge to the current nearest neighbor. The inner loop evaluates all remaining points in $S$, determining whether to prune them based on the pruning rule. In the non-truncated SNG pruning process, the outer loop iterates at most $O(n^{2/3+\epsilon})$ times, as established by Theorem 1. The inner loop evaluates all points in the candidate set $S$, excluding the nearest neighbor, checking up to n2n-2 points per iteration. This yields a complexity upper bound of $O(n^{5/3+\epsilon})$, tighter than the $O(n^{2})$ estimate in [30]. For the truncated variant, RobustPrune, the outer loop runs at most $R$ times, where $R$ is the optimized truncation parameter. The inner loop evaluates up to n2n-2 points, resulting in a worst-case complexity of $O(R\cdot n)$.
@@
Total Complexity of Vamana. The following are the graph construction steps of the Vamana algorithm used in DiskANN. To begin with, the graph is randomly initialized as an $R$-regular graph. $R$ is the maximum degree specified by the parameter sweep. To reduce the indexing time, for each point $p$ in the dataset, GreedySearch is first performed starting from the centroid $s$ to obtain the set of volunteer points, which are then used as candidate points for pruning, the complexity can be approximate by $O(R\cdot logn)$ and written as $C_{1}\cdot R\log n+b_{1}$ [12]. Next, we apply RobustPrune to prune the randomized graph. This process runs for at most $R$ iterations, and in each iteration, the number of points to be checked is approximately $O(R\log n)$. Considering the parameter $\alpha$, as $\alpha$ increases, the graph has higher degree, hence the complexity can be regarded as inversely proportional to $\alpha$. Therefore, the algorithm’s complexity is estimated to be $O(R\cdot R\log n/\alpha)$ and can be expressed as $C_{2}(R\cdot R\log n/\alpha)+b_{2}$. To enhance GreedySearch convergence during the search phase, we add reverse edges for each neighbor of point $p$. This may cause a degree overrun beyond $R$. In such cases, RobustPrune is reapplied. For each point $p$ with at most $R$ possible neighbors, we add reverse edges and reconstruct the graph. The candidate set size does not exceed $R+1$, and the inner loop iterates at most $R$ times. A larger $\alpha$ increases the number of neighbors, raising the probability of a degree overrun, which we model as a proportional relationship, so the complexity is expressed as $C_{3}(\alpha\cdot R\cdot R^{2})+b_{3}$.
Appendix G Discussion
@@
As elaborated in [26], a widely used model is the Poisson point process, which assumes that the number of points in any bounded region follows a Poisson distribution and that point locations are independent and identically distributed. Crucially, given a fixed number of points within a bounded region, the conditional distribution of those points under a Poisson point process is uniform. This property provides a natural justification for analyzing nearest neighbor graph constructions under the uniform distribution assumption, as the conditioning step aligns with the typical setup in ANNS frameworks such as DiskANN and SNG.
@@
Appendix H Extended Results
H.1 Informations of Datasets
Below are the detailed specifications of the three datasets used, including their size, the size of the query set, and their dimension.
Table 2: Datasets
Dataset NB Base Vectors NB Query Vectors Dimension
SIFT1M 1,000,000 10,000 128
GIST1M 1,000,000 1,000 960
DEEP1M 1,000,000 1,000 256
H.2 Pruning process on SIFT1M
Figure 8: Left Figure (General Pruning Process): The pruning process for a sample point from SIFT1M is illustrated in the left figure. The horizontal axis represents the number of iteration steps, while the vertical axis indicates the number of processed points. Notably, only 4 steps are required to exceed the 80% threshold. However, the remaining pruning iterations are significantly prolonged, with the final 20% of points requiring 13 times the iterations of the initial points. This indicates that the algorithm’s speed diminishes quickly, leading to a slower rate of subsequent point scanning, with later stages consuming the majority of the scanning time. Right Figure (An Extreme Case): The pruning process for the mean point shows a similar trend of rapid initial progress followed by a slow phase. However, the slower pruning phase dominates, requiring 1092 steps to complete.
H.3 Query performance on GMM
Below are the test results for the search performance on the GMM dataset, where latency is measured in microseconds. It can be observed that the latency corresponding to a 99% recall rate does not exceed 3 ms.
Table 3: Query Processing Performance on GMM
L / Beamwidth QPS Mean Latency 99.9 Latency Mean IO (us) CPU (s) Recall@10
10 / 2 35,187.99 768.96 2,115.00 746.02 3.90 97.23
20 / 2 19,995.63 1,442.69 2,676.00 1,403.40 6.15 99.96
30 / 2 13,505.18 2,048.59 3,706.00 1,992.13 10.46 100.00
40 / 2 10,163.05 2,723.53 4,740.00 2,652.44 11.80 100.00
50 / 2 8,135.30 3,375.69 5,658.00 3,289.35 14.62 100.00
100 / 2 4,058.43 6,859.58 24,449.00 6,695.53 29.03 100.00
H.4 Degree Distribution of GMM for other $\alpha$ values
Below, we present the degree distribution of the GMM dataset under different $\alpha$ values, all of which conform to the theoretical bound.
Figure 9: Degree distribution of GMM when $\alpha=1.00$
Figure 10: Degree distribution of GMM when $\alpha=1.25$
Figure 11: Degree distribution of GMM when $\alpha=1.50$

### v3

@@ Appendix C Complexity Analysis Details
Below is a detailed complexity analysis of each component in the SNG construction algorithm. Complexity of GreedySearch. The GreedySearch algorithm iteratively selects the closest neighbor to the query and follows a path to an approximate nearest neighbor, evaluating all neighbors at each step. According to (NSG19), the complexity of GreedySearch is determined by the product of the average path length and the average degree. Theorem 1 establishes that the path length is $O(\log n)$. Thus, for a graph with degree $K$, the search complexity is $O(K\cdot\log n)$.
@@
Complexity of Pruning. In the SNG pruning algorithm, the outer loop checks if the candidate set $S$ is empty. If not, it adds an edge to the current nearest neighbor. The inner loop evaluates all remaining points in $S$, determining whether to prune them based on the pruning rule. In the non-truncated SNG pruning process, the outer loop iterates at most $O(n^{2/3+\epsilon})$ times, as established by Theorem 4. The inner loop evaluates all points in the candidate set $S$, excluding the nearest neighbor, checking up to n2n-2 points per iteration. This yields a complexity upper bound of $O(n^{5/3+\epsilon})$, tighter than the $O(n^{2})$ estimate in (NEURIPS2019). For the truncated variant, SNG-Prune, the outer loop runs at most $R$ times. The inner loop evaluates up to n2n-2 points, resulting in a worst-case complexity of $O(R\cdot n)$.
@@
Total Complexity of Construction od SNG. To begin with, the graph is randomly initialized as an $R$-regular graph. $R$ is the maximum degree specified by the parameter sweep. To reduce the indexing time, for each point $p$ in the dataset, GreedySearch is first performed starting from the centroid $s$ to obtain the set of volunteer points, which are then used as candidate points for pruning, the complexity can be approximate by $O(R\cdot logn)$ and written as $C_{1}\cdot R\log n+b_{1}$ (NSG19). Next, we apply SNG-tPrune to prune the randomized graph. This process runs for at most $R$ iterations, and in each iteration, the number of points to be checked is approximately $O(R\log n)$. Considering the parameter $\alpha$, as $\alpha$ increases, the graph has higher degree, hence the complexity can be regarded as inversely proportional to $\alpha$. Therefore, the algorithm’s complexity is estimated to be $O(R\cdot R\log n/\alpha)$ and can be expressed as $C_{2}(R\cdot R\log n/\alpha)+b_{2}$. To enhance GreedySearch convergence during the search phase, we add reverse edges for each neighbor of point $p$. This may cause a degree overrun beyond $R$. In such cases, SNG-tPrune is reapplied. For each point $p$ with at most $R$ possible neighbors, we add reverse edges and reconstruct the graph. The candidate set size does not exceed $R+1$, and the inner loop iterates at most $R$ times. A larger $\alpha$ increases the number of neighbors, raising the probability of a degree overrun, which we model as a proportional relationship, so the complexity is expressed as $C_{3}(\alpha\cdot R\cdot R^{2})+b_{3}$.
Appendix D Discussion
@@
As elaborated in (ICML2020), a widely used model is the Poisson point process, which assumes that the number of points in any bounded region follows a Poisson distribution and that point locations are independent and identically distributed. Crucially, given a fixed number of points within a bounded region, the conditional distribution of those points under a Poisson point process is uniform. This property provides a natural justification for analyzing nearest neighbor graph constructions under the uniform distribution assumption, as the conditioning step aligns with the typical setup in ANNS frameworks such as DiskANN and SNG.